Massachusetts Institute of Technology Handout 4 6.042J/18.062J: Mathematics for Computer Science September 10, 1998 Professors Nancy Lynch and Srinivas Devadas

Schedules

Tutorial Schedule on Fridays

Time  Room
R01: 10-12 8-119 Nancy Lynch/Srini Devadas
R02: 10-12 2-142  Rodrigo Leroux
R03: 11-1 26-210  Brian Dean
R04: 1-3 26-168  Brad Banks
R05: 1-3 36-839  Rodrigo Leroux
R06: 2-4 26-310  Brian Dean
R07: 2-4 24-402  Sarfraz Khurshid
R08: 10-12 36-372  Sarfraz Khurshid
R09: 2-4 24-407  Adam Gould
R10: 2-4 36-144  John Dunagan
R11: 9-11 26-168  Alex Duran
R12: 3-5 26-168  Brad Banks
R13: 11-1 2-139  Alex Duran
R16: 11-1 2-147  Adam Gould

2 Handout 4: Schedules

Office Hours

Time  Room
Nancy Lynch  Wednesdays 4-5 NE43-365 x3-7225
Srini Devadas  Mondays 1-3 NE43-258 x3-0454
Brad Banks  Mondays 1-3  24-311 x8-5690
Brian Dean  Mondays 11-1  24-311 x8-5690
Alex Duran  Mondays 5-8  24-311 x8-5690
Adam Gould  Tuesdays, Thursdays 12:30-2:00  24-311 x8-5690
Sarfraz Khurshid  Wednesdays 3-5  24-311 x8-5690
Eric Lehman  Mondays 2-5  24-311
Rodrigo Leroux  Tuesdays 10-12  24-311 x8-5690
John Dunagan  Mondays 9-11 NE43-311 x3-2345